A 300 ml water sample was taken from a stream in Mexico City and analyzed for dissolved oxygen. Using the winkler method, 58.88 ml of 2.42 x 10^-3 M sodium thiosulfate titrant was required. Calculate the dissolved oxygen in the water and express this as mg/L.


1
Expert's answer
2021-01-19T04:40:30-0500

2S2O"_3^{-2}" + I"_3^-" "\to" 3I"^-" + S4O"_6^{-2}"

1000ml of 1N Na2S2O3 = 8g dissolved oxygen

1N Na2S2O3 = (1×2) M Na2S2O3

1000ml of 2M Na2S2O3 = 8g dissolved oxygen

2000 mmol Na2S2O3 = 8g

1 m mol Na2S2O3 = "\frac{8}{2000}" g

55.88 ×2.42×10-3 m mol Na2S2O3 = "\frac{8×55.88×2.42×10^-³}{2000}"

= 0.0005409184 mg/L